A soil sample was taken with a cylindrical metal core having a volume of 300cm3. the soil was saturated with water and then drainage water was allowed to drip. Meanwhile, the evaporation of water from the soil was prevented. The drained soil sample weighed 500g before oven-drying and 350g after oven-drying. Calculate the following:
Bulk Density
MC% by weight and volume
Porosity(assume Dp=2.65g/cm3)
cm of water per 100cm depth of soil (depth of water in cm = MC% by volume x depth of soil) Show calculations
1
Expert's answer
2021-10-15T12:32:31-0400

Volume of metal core V = 300 cm3

Saturated soil weight "w_s = 500 \;g"

Dry soil weight "w_d = 350 \;g"

Bulk Density "V_b = \frac{w_s}{V} = \frac{500}{300} = 1.67 \;g/cm^3"

MC% by weight and volume

"MC_{w} = \frac{w_s-w_d}{w_d} \times 100 \\

= \frac{500-350}{350} \times 100 \\

= 42.85 \;\%"

Dry density

"ρ_d = \frac{w_d}{V} \\

= \frac{350}{300}=1.167 \;g/cm^3 \\

ρ_d = \frac{D_pρ_w }{1 +e} \\

1.167 = \frac{2.65 \times 1}{1+e} \\

e = 1.27"

Porosity

"η = \frac{e}{1+e} = 0.559 \\

\frac{Volume \; of \; water}{Total \; volume} = \frac{V_w}{V} = 0.559"

Since soil is 100% saturated with water.

"MC_{v} = \frac{V_w}{V} \times 100 \\

= 0.559 \times 100 \\

= 55.9 \;\%"

cm of water per 100cm depth of soil

depth of water in cm = MC % by volume "\times" depth of soil

= 55.9 % "\times" 100 cm

= 55.9 cm